# Can AI-Generated Persuasion Be Detected? Persuaficial Benchmark and AI vs. Human Linguistic Differences

### v1

@@ Abstract
Arkadiusz Modzelewski1,2,3, Paweł Golik, Anna Kołos3, Giovanni Da San Martino1 1University of Padua, Italy 2Polish-Japanese Academy of Information Technology, Poland 3NASK National Research Institute, Poland Correspondence: contact@amodzelewski.com
@@ 1 Introduction
To address RQ2, we conducted an analysis using the StyloMetrix tool [^1] https://github.com/ZILiAT-NASK/StyloMetrix, which generates fully interpretable and reproducible vectors representing a wide range of linguistic features in text Okulska et al. (2023). Prior work shows that StyloMetrix performs well on persuasion detection with classical machine learning Modzelewski et al. (2023, 2024), underscoring its suitability for studying the linguistic features of persuasive texts. In our analysis, we examined the full range of linguistic features offered by open-source StyloMetrix.
@@
We release our codebase, dataset and all prompts222anonymous.4open.science/r/Persuaficial-26EC/.
@@ 2 Human Persuasion Datasets
  • SemEval 2023 Task 3 Dataset: A multilingual, multifaceted collection of online news articles annotated with various persuasion techniques on paragraph level Piskorski et al. (2023d). Its taxonomy and dataset are widely adopted within the NLP community for persuasion research Barrón-Cedeño et al. (2024); Dimitrov et al. (2024); Modzelewski et al. (2025). This dataset was introduced as part of SemEval 2023 Task 3 on persuasion detection Piskorski et al. (2023b).
@@
  • ChangeMyView: A dataset derived from the Reddit ChangeMyView discussion community. Dataset contains 3,051 conversations in which the persuader tries to convince the persuadee to change their mind Tan et al. (2016). It is commonly used in persuasion research Ziems et al. (2024); Wei et al. (2016); Dutta et al. (2020).
@@ 3.3 Persuaficial Quality Evaluation
As mentioned, for the Open-ended generation setting, we first summarize each selected persuasive text into a short <sentence>. To ensure these <sentence>s accurately represent the source human text, we conducted a human evaluation following explicit and rigorous annotation guidelines (see Appendix D.1).
@@ 5 Linguistic Differences Between Machine and Human Persuasion
In this section, we investigate the linguistic differences between human-written and AI-generated persuasive texts. We focus on English due to the limited availability of high-quality datasets in other languages. While the SemEval 2023 Task 3 data provides a multilingual resource Srba et al. (2024); Piskorski et al. (2023c), relying solely on it could introduce dataset-specific biases. To mitigate this, we use English part of Persuaficial and human-written counterparts from three well-established datasets.
@@ 5.1 Our Approach for Linguistic Analysis
To investigate the linguistic differences between human-written persuasive texts and LLM-generated persuasive texts, we adopt an explainable, feature-based analysis grounded in stylometry. Our objective is to identify the linguistic features that most strongly differentiate LLM-generated persuasive texts from human-written ones. For each linguistic feature, we compare the distributions of the two groups using effect-size-based analysis together with significance testing. Effect sizes quantify the magnitude of the difference between human and AI-generated texts for each feature Frey (2021), while significance testing evaluates whether these distributional differences are statistically meaningful. Our analysis identifies which linguistic properties differ systematically between human- and AI-produced persuasive texts.
@@ 5.2 Experimental Setup
We first represent each persuasive human text and its AI-generated counterpart using StyloMetrix Okulska et al. (2023). For each text, we calculate a 196-dimensional vector of linguistic features. This results in a tabular representation, where each row corresponds to a text encoded by its computed linguistic features. We utilize StyloMetrix, because it is an open-source tool that provides fully interpretable, linguistically grounded feature representations. Moreover, prior work has demonstrated its effectiveness for persuasion detection using classical machine learning models Modzelewski et al. (2023, 2024), confirming its suitability for analyzing the linguistic characteristics of persuasive texts. Finally, to further justify our choice, we show that StyloMetrix features with classical machine learning can distinguish human-written from AI-generated persuasive texts in Appendix H.
For each linguistic feature, we directly compare its distribution in human-written and LLM-generated persuasive texts, conducting this analysis separately for each generation approach. We utilize Cohen’s d statistic (see Appendix I for definiton and how we computed it in our experiments) which is a type of effect size measure used to represent the magnitude of differences between two groups on a given variable Frey (2021). To evaluate whether feature distributions significantly differ between human and AI-generated texts, we perform Wilcoxon signed-rank tests Wilcoxon (1945), a non-parametric paired test appropriate for comparing matched text pairs Peyrard et al. (2021); Dror et al. (2018). Wilcoxon signed-rank tests have seen widespread adoption in the NLP community Karwa and Singh (2025); Zhou et al. (2025); Ciaccio et al. (2025), including in studies comparing feature distributions Stodden and Kallmeyer (2020).
@@
Figure 1: StyloMetrix features with the highest discriminative importance in distinguishing human-written from LLM-generated persuasive text.
@@ 5.3 Results and Analysis
We computed Cohen’s d values for 196 linguistic features across four generation approaches and four LLMs utilized to generate texts for our Persuaficial dataset. Table 5 sorts top linguistic features by the absolute Cohen’s d ($|C_{d}|$) values per model and generation approach ($G_{P}$). Our analysis and discussion is based on the twenty features with the largest $|C_{d}|$ for each model and generation strategy (more detailed tables available in Appendix J). Statistical analysis using Wilcoxon signed-rank tests confirmed that all twenty features for each scenario exhibit significant distributional differences between human-written and AI-generated persuasive texts. Figure 1 provides definitions of the key differentiating features.
High values of features such as L_CONT_T (the proportion of unique content-word forms relative to total tokens), LTOKEN_RATIO_LEM (the ratio of unique lemmas to total tokens), and L_CONT_A (the proportion of tokens that are content words) indicate that AI-generated texts tend to contain more varied content words and higher informational density per sentence. These patterns suggest that lexical diversity and content richness are characteristic markers of AI authorship. Similarly, low values for ST_REPET_WORDS are indicative of AI-generated persuasive texts, suggesting that reduced word repetition serves as a strong signal of LLM-generated text. A higher proportion of function words (L_FUNC_A) indicates that a persuasive text is likely human-written. This means that frequent use of grammatical connectors (such as articles, prepositions, pronouns, and auxiliary verbs) is a signal of human text. Furthermore, AI-generated persuasive texts generally exhibit lower punctuation density, especially in texts from Llama and GPT-4.1-mini. However, certain punctuation marks, including commas (L_PUNCT_COM) and dashes (L_PUNCT_DASH), occur more frequently in these AI texts. The rarity of syntactically marked constructions, such as inversions (SY_INV_PATTERNS), is a distinguishing feature of AI text, as these complex syntactic patterns are more typical of human-written persuasive texts.
@@
Feature Name $C_{d}$ Feature Name $C_{d}$ Feature Name $C_{d}$ Feature Name $C_{d}$
@@
L_CONT_T 0.51 L_CONT_T 0.70 L_CONT_T 0.77 L_CONT_T 1.62
L_CONT_A 0.46 L_CONT_A 0.62 L_CONT_A 0.77 L_CONT_A 1.43
SY_INV_PATTERNS -0.45 VT_MIGHT 0.62 L_PUNCT_DASH 0.72 LTOKEN_RATIO_LEM 1.32
LTOKEN_RATIO_LEM 0.37 SY_INV_PATTERNS -0.61 L_FUNC_A -0.57 SENT_D_NP 1.05
L_FUNC_A -0.36 L_PLURAL_NOUNS 0.54 LTOKEN_RATIO_LEM 0.51 ST_REPET_WORDS -1.00
@@
SENT_ST_WPERSENT 0.71 SY_INV_PATTERNS -0.75 SENT_ST_WPERSENT 0.80 VF_INFINITIVE 1.34
SENT_ST_DIFFER -0.70 G_ACTIVE -0.72 L_ADJ_POSITIVE 0.72 SY_IMPERATIVE 1.24
L_PUNCT_COM 0.67 SENT_ST_WPERSENT 0.70 L_PUNCT_COM 0.71 G_ACTIVE -0.91
SY_INV_PATTERNS -0.61 L_CONT_T 0.62 L_CONT_T 0.66 SENT_D_VP 0.83
L_CONT_T 0.54 SENT_D_PP 0.62 POS_ADJ 0.65 L_PUNCT_DOT -0.81
@@
L_CONT_T 0.75 L_CONT_T 1.02 L_CONT_T 1.04 SY_IMPERATIVE 1.74
L_CONT_A 0.67 L_CONT_A 1.02 L_CONT_A 1.01 L_CONT_T 1.4
L_FUNC_A -0.62 L_FUNC_A -0.89 L_FUNC_A -0.97 L_CONT_A 1.25
SY_INV_PATTERNS -0.61 POS_PRO -0.73 L_ADJ_POSITIVE 0.90 SENT_D_NP 1.16
L_ADV_SUPERLATIVE 0.43 SY_INV_PATTERNS -0.71 POS_ADJ 0.83 PS_CAUSE -1.15
@@
L_CONT_A 0.73 L_CONT_A 0.92 L_CONT_A 1.02 L_CONT_T 1.57
L_CONT_T 0.71 L_CONT_T 0.92 L_CONT_T 0.99 L_CONT_A 1.37
SY_INV_PATTERNS -0.58 L_FUNC_A -0.73 L_FUNC_A -0.82 SY_IMPERATIVE 1.32
L_FUNC_A -0.54 SY_INV_PATTERNS -0.72 L_ADJ_POSITIVE 0.78 LTOKEN_RATIO_LEM 1.20
L_PUNCT_COM 0.49 POS_NOUN 0.59 POS_ADJ 0.72 SY_EXCLAMATION 1.11
@@ 6 Related Work
Parallel work has investigated approaches for detecting persuasion. Da San Martino et al. (2019) presented corpus of news annotated at the fragment level with 18 persuasive techniques and proposed multi-granularity neural network to detect persuasion. Piskorski et al. (2023d) extended the taxonomy of techniques proposed by Da San Martino et al. (2019) and presented a multilingual dataset. Moreover, persuasion detection was a task of different recognized workshops such SemEval or SlavicNLP Da San Martino et al. (2020); Dimitrov et al. (2021); Piskorski et al. (2023b, 2025).
@@ 7 Conclusion
Through a detailed analysis, we identify key linguistic differences between human and AI-generated persuasive texts. AI-generated texts tend to exhibit higher lexical diversity, increased content-word density, and lower function-word usage, while complex syntactic patterns are more characteristic of human persuasive writing. Text generation approaches further modulate these features: intensified persuasion amplifies adverbial modifiers, whereas subtle persuasion relies on modal hedges and declarative constructions.
@@ Appendix A Human Dataset - Rationale Behind Our Choice
DIPROMATS 2024 (Moral et al., 2024), which build on the SemEval task, offers data that is directly compatible with this definition and is therefore suitable for our experiments. Finally, Piskorski et al. (2025) demonstrate that this definition can be effectively and reliably applied to debates and discussions, motivating our choice of the ChangeMyView dataset, comprising message exchanges between persuaders and persuadees, as an additional human-written source.
@@ G.2 Detailed results on non-English texts of Persuaficial dataset
Appendix I Cohen’s d statistic definition and our computation
For each linguistic feature, we calculate Cohen’s $d$ to quantify the magnitude of the shift between the feature distribution of generated texts $g_{i}$ and that of their human-written persuasive counterparts $r_{i}$. Cohen’s $d$ is defined in Equation 5 and is computed as follows.
First, we calculate the mean of each feature for human-written and generated texts:
$\bar{r}=\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}r_{i},\qquad\bar{g}=\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}g_{i},$ (1)
where $n_{r}$ and $n_{g}$ denote the number of human-written and generated texts, respectively. In our experiments, $n_{r}=n_{g}$ as for each human-written persuasive text we have AI-generated counterpart.
Next, we compute the sample standard deviations for each group:
$s_{r}=\sqrt{\frac{1}{n_{r}-1}\sum_{i=1}^{n_{r}}(r_{i}-\bar{r})^{2}},$ (2)
$s_{g}=\sqrt{\frac{1}{n_{g}-1}\sum_{i=1}^{n_{g}}(g_{i}-\bar{g})^{2}}.$ (3)
Using these, we calculate the pooled standard deviation:
$s_{\text{pooled}}=\sqrt{\frac{(n_{r}-1)s_{r}^{2}+(n_{g}-1)s_{g}^{2}}{n_{r}+n_{g}-2}}.$ (4)
Finally, Cohen’s $d$ is obtained as the difference in means normalized by the pooled standard deviation:
$d=\frac{\bar{g}-\bar{r}}{s_{\text{pooled}}}$ (5)
This effect size, Cohen’s $d$, provides a standardized measure of the shift in feature distributions between generated and human-written persuasive texts, allowing comparison across features and models.
Since many feature distributions may be non-Gaussian, we avoid this assumptions when testing for a significance of the shift. Statistical significance is assessed with a paired Wilcoxon signed-rank test on the per text differences $d_{i}=g_{i}-r_{i}$, with Benjamini-Hochberg FDR correction across features. As we perform one test per feature, we control for multiple comparisons using the Benjamini-Hochberg false discovery rate (FDR) correction Benjamini and Hochberg (1995) across all features. We report a significance indicator in our tables.

### v2

@@ Abstract
Arkadiusz Modzelewski1,2,3, Paweł Golik, Anna Kołos1, Giovanni Da San Martino2 1NASK National Research Institute, Poland 2 University of Padua, Italy 3 Polish-Japanese Academy of Information Technology, Poland Correspondence: contact@amodzelewski.com
@@ 1 Introduction
To address RQ2, we conducted an analysis using the StyloMetrix tool [^1] https://github.com/ZILiAT-NASK/StyloMetrix, which generates fully interpretable and reproducible vectors representing a wide range of linguistic features in text Okulska et al. (2023). Prior work shows that StyloMetrix performs well on persuasion detection with classical machine learning Modzelewski et al. (2023, 2024b), underscoring its suitability for studying the linguistic features of persuasive texts. In our analysis, we examined the full range of linguistic features offered by open-source StyloMetrix.
@@
We release our codebase, dataset and all prompts [^2] https://github.com/ArkadiusDS/Persuaficial.
@@ 2 Human Persuasion Datasets
  • SemEval 2023 Task 3 Dataset: A multilingual, multifaceted collection of online news articles annotated with various persuasion techniques on paragraph level Piskorski et al. (2023d). Its taxonomy and dataset are widely adopted within the NLP community for persuasion research Barrón-Cedeño et al. (2024); Dimitrov et al. (2024); Modzelewski et al. (2025). This dataset was introduced as part of SemEval 2023 Task 3 on persuasion detection Piskorski et al. (2023b). The extended taxonomy was also applied to annotate a parliamentary debates and social media dataset presented at the SlavicNLP workshop Piskorski et al. (2025a).
@@
  • ChangeMyView: A dataset derived from the Reddit ChangeMyView discussion community. Dataset contains 3,051 conversations in which the persuader tries to convince the persuadee to change their mind Tan et al. (2016). It is commonly used in persuasion research Ziems et al. (2024); Wei et al. (2016); Dutta et al. (2020); Labruna et al. (2026).
@@ 3.3 Persuaficial Quality Evaluation
As mentioned, for the Open-ended generation setting, we first summarize each selected persuasive text into a short <sentence>. To ensure these LLM-generated <sentence>s accurately reflect the content of the corresponding human texts, we conducted a human evaluation following explicit and rigorous annotation guidelines (see Appendix D.1).
@@ 5 Linguistic Differences Between Machine and Human Persuasion
In this section, we investigate the linguistic differences between human-written and AI-generated persuasive texts. We focus on English due to the limited availability of high-quality datasets annotated with persuasion in other languages. While the SemEval 2023 Task 3 data provides a multilingual resource Srba et al. (2024); Piskorski et al. (2023c), relying solely on it could introduce dataset-specific biases. To mitigate this, we use English part of Persuaficial and human-written counterparts from three well-established datasets.
In our analysis, we focus on linguistic features that encompass morphosyntactic dimensions (e.g., inflection, POS distribution), syntax, lexis, and punctuation, among others.
@@ 5.1 Our Approach for Linguistic Analysis
To investigate the linguistic differences between human-written persuasive texts and LLM-generated persuasive texts, we adopt an explainable, feature-based analysis grounded in stylometry. Our objective is to identify the linguistic features that most strongly differentiate LLM-generated persuasive texts from human-written ones. For each linguistic feature, we compare the distributions of the two groups using effect-size-based analysis together with significance testing. Effect sizes quantify the magnitude of the difference between human and AI-generated texts for each feature Frey (2021), while significance testing evaluates whether these distributional differences are statistically meaningful. We restrict our analysis to persuasive texts and aim to characterize differences between human and AI-generated persuasive text, without attempting to isolate persuasion-specific features from broader AI-human writing differences.
@@ 5.2 Experimental Setup
We first represent each persuasive human text and its AI-generated counterpart using StyloMetrix Okulska et al. (2023). For each text, we calculate a 196-dimensional vector of linguistic features. This results in a tabular representation, where each row corresponds to a text encoded by its computed linguistic features. We utilize StyloMetrix, because it is an open-source tool that provides fully interpretable, linguistically grounded feature representations. Moreover, prior work has demonstrated its effectiveness for persuasion detection using classical machine learning models Modzelewski et al. (2023, 2024b), confirming its suitability for analyzing the linguistic characteristics of persuasive texts. Finally, to further justify our choice, we show that StyloMetrix features with classical machine learning can distinguish human-written from AI-generated persuasive texts in Appendix H.
For each linguistic feature $j$, we directly compare its distribution in human-written and LLM-generated persuasive texts, conducting this analysis separately for each generation approach. We utilize paired Cohen’s $d$ statistic which is a type of effect size measure used to represent the magnitude of differences between two groups on a given variable Frey (2021). We report the paired version of Cohen’s $d$, as our analysis is based on matched pairs consisting of AI-generated persuasive texts and their corresponding human-written counterparts.
In our calculation, we firstly compute the element-wise difference between the generated and real feature values across all $n=1,500$ persuasive text pairs (500 per dataset, across three datasets):
$\Delta_{j}^{i}=g_{j}^{i}-r_{j}^{i},\quad i=1,\ldots,1500$ (1)
where $g_{j}^{i}$ denotes the value of feature $j$ for the $i$-th AI-generated text and $r_{j}^{i}$ denotes that of the corresponding human-written text. To measure the magnitude of the shift introduced by the language model, we compute paired Cohen’s $d$:
$d_{j}=\frac{\bar{\Delta}_{j}}{s_{\Delta_{j}}}$ (2)
where
$\bar{\Delta}_{j}=\frac{1}{n}\sum_{i=1}^{n}\Delta_{j}^{i}$ (3)
is the mean of the paired differences for feature $j$, and
$s_{\Delta_{j}}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(\Delta_{j}^{i}-\bar{\Delta}_{j})^{2}}$ (4)
is the standard deviation of the paired differences. This procedure is repeated independently for each feature $j$, each generating model, and each generation technique.
A positive Cohen’s $d$ indicates the AI-generated text exhibits higher values of a given feature compared to the human original, while a negative value indicates a decrease.
To evaluate whether feature distributions significantly differ between human and AI-generated texts, we perform Wilcoxon signed-rank tests Wilcoxon (1945), a non-parametric paired test appropriate for comparing matched text pairs Peyrard et al. (2021); Dror et al. (2018). Wilcoxon signed-rank tests have seen widespread adoption in the NLP community Karwa and Singh (2025); Zhou et al. (2025); Ciaccio et al. (2025), including in studies comparing feature distributions Stodden and Kallmeyer (2020). See Appendix I for more details regarding our statistical analysis.
@@
Figure 1: StyloMetrix features that exhibit some of the strongest discriminative power for distinguishing human-written persuasive text from persuasive content generated by LLMs.
@@ 5.3 Results and Analysis
We computed Cohen’s d values for 196 linguistic features across four generation approaches and four LLMs utilized to generate texts for our Persuaficial dataset. Table 5 sorts top linguistic features by the absolute Cohen’s d ($|d_{j}|$) values per model and generation approach ($G_{P}$). Our analysis and discussion is based on the twenty features with the largest $|d_{j}|$ for each model and generation strategy (more detailed tables available in Appendix J). Statistical analysis using Wilcoxon signed-rank tests confirmed that all twenty features for each scenario exhibit significant distributional differences between human-written and AI-generated persuasive texts. Figure 1 provides definitions of the key differentiating features.
High values of features such as L_CONT_T (the proportion of unique content-word forms relative to total tokens), LTOKEN_RATIO_LEM (the ratio of unique lemmas to total tokens), and L_CONT_A (the proportion of tokens that are content words) indicate that AI-generated texts tend to contain more varied content words and higher informational density per sentence. These patterns suggest that lexical diversity and content richness are characteristic markers of AI authorship. Similarly, low values for ST_REPET_WORDS are indicative of AI-generated persuasive texts, suggesting that reduced word repetition serves as a strong signal of LLM-generated text. A higher proportion of function words (L_FUNC_A) indicates that a persuasive text is likely human-written. This means that frequent use of grammatical connectors (such as articles, prepositions, pronouns, and auxiliary verbs) is a signal of human text. Furthermore, AI persuasive texts generated by Llama tend to have a lower density of periods. However, certain punctuation marks, including commas (L_PUNCT_COM) and dashes (L_PUNCT_DASH), occur more frequently in AI texts. The rarity of syntactically marked constructions, such as inversions (SY_INV_PATTERNS), is a distinguishing feature of AI text, as these complex syntactic patterns are more typical of human-written persuasive texts.
@@
Feature Name $d_{j}$ Feature Name $d_{j}$ Feature Name $d_{j}$ Feature Name $d_{j}$
@@
LTOKEN_RATIO_LEM 0.92 LTOKEN_RATIO_LEM 0.94 L_CONT_T 1.18 L_CONT_T 1.41
L_CONT_T 0.82 L_CONT_T 0.82 LTOKEN_RATIO_LEM 1.09 L_CONT_A 1.23
ST_REPET_WORDS -0.75 ST_REPET_WORDS -0.76 L_CONT_A 1.02 LTOKEN_RATIO_LEM 1.08
L_CONT_A 0.65 L_PLURAL_NOUNS 0.66 L_FUNC_A -0.86 ST_REPET_WORDS -0.85
L_FUNC_A -0.58 L_CONT_A 0.63 ST_REPET_WORDS -0.82 SENT_D_NP 0.76
@@
SENT_ST_WPERSENT 0.77 SENT_ST_WPERSENT 0.73 L_CONT_T 0.95 VF_INFINITIVE 1.01
SENT_ST_DIFFER -0.77 L_CONT_T 0.70 SENT_ST_WPERSENT 0.88 SY_IMPERATIVE 0.89
L_CONT_T 0.76 SENT_D_PP 0.67 L_CONT_A 0.76 G_ACTIVE -0.70
L_PUNCT_COM 0.68 SENT_ST_DIFFER -0.64 L_ADJ_POSITIVE 0.75 SENT_D_VP 0.68
LTOKEN_RATIO_LEM 0.61 L_PROPER_NAME -0.61 SENT_ST_DIFFER -0.73 L_OUR_PRON 0.60
@@
L_CONT_T 0.87 L_CONT_T 1.02 L_CONT_T 1.11 SY_IMPERATIVE 1.25
LTOKEN_RATIO_LEM 0.73 L_CONT_A 0.87 L_FUNC_A -0.95 L_CONT_T 1.21
L_FUNC_A -0.70 LTOKEN_RATIO_LEM 0.85 L_CONT_A 0.95 L_CONT_A 1.05
L_CONT_A 0.68 L_FUNC_A -0.76 L_ADJ_POSITIVE 0.83 L_FUNC_A -0.92
ST_REPET_WORDS -0.67 ST_REPET_WORDS -0.67 POS_ADJ 0.77 PS_CAUSE -0.89
@@
L_CONT_T 0.90 L_CONT_T 1.02 L_CONT_T 1.23 L_CONT_T 1.34
LTOKEN_RATIO_LEM 0.82 L_CONT_A 0.87 L_CONT_A 1.12 L_CONT_A 1.16
L_CONT_A 0.79 LTOKEN_RATIO_LEM 0.85 L_FUNC_A -0.98 LTOKEN_RATIO_LEM 0.97
ST_REPET_WORDS -0.72 L_FUNC_A -0.76 LTOKEN_RATIO_LEM 0.89 SY_IMPERATIVE 0.93
L_FUNC_A -0.66 ST_REPET_WORDS -0.67 L_ADJ_POSITIVE 0.87 ST_REPET_WORDS -0.83
@@ 6 Related Work
Parallel work has investigated approaches for detecting persuasion. Da San Martino et al. (2019) presented corpus of news annotated at the fragment level with eighteen persuasive techniques and proposed multi-granularity neural network to detect persuasion. Piskorski et al. (2023d) extended the taxonomy of techniques proposed by Da San Martino et al. (2019) and presented a multilingual dataset. Modzelewski et al. (2024a) introduced a Polish disinformation dataset annotated with eleven manipulation techniques, defining manipulation as the intentional use of persuasion for malicious purposes. Moreover, persuasion detection was a task of different recognized workshops such SemEval or SlavicNLP Da San Martino et al. (2020); Dimitrov et al. (2021); Piskorski et al. (2023b, 2025b).
@@ 7 Conclusion
In addition, through a detailed analysis, we identify key linguistic differences between human and AI-generated persuasive texts. AI-generated texts tend to exhibit higher lexical diversity, increased content-word density, and lower function-word usage, while complex syntactic patterns are more characteristic of human persuasive writing. Text generation approaches further modulate these features: intensified persuasion amplifies adverbial modifiers, whereas subtle persuasion relies on modal hedges and declarative constructions.
@@
Acknowledgments
Giovanni Da San Martino would like to thank the Qatar National Research Fund, part of Qatar Research Development and Innovation Council (QRDI), for funding this work by grant NPRP14C0916-210015.
@@ Appendix A Human Dataset - Rationale Behind Our Choice
DIPROMATS 2024 (Moral et al., 2024), which build on the SemEval task, offers data that is directly compatible with this definition and is therefore suitable for our experiments. Finally, Piskorski et al. (2025b) demonstrate that this definition can be effectively and reliably applied to debates and discussions, motivating our choice of the ChangeMyView dataset, comprising message exchanges between persuaders and persuadees, as an additional human-written source.
@@ G.2 Detailed results on non-English texts of Persuaficial dataset
Appendix I Statistical Analysis with Wilcoxon Signed-rank Test
To assess whether the observed shifts are statistically significant without assuming normality of the paired differences, we apply the Wilcoxon signed-rank test:
$H_{0}:\tilde{\mu}_{\Delta}=0\quad\text{vs.}\quad H_{1}:\tilde{\mu}_{\Delta}\neq 0$ (5)
where $\tilde{\mu}_{\Delta}=\text{median}(\Delta_{j}^{1},\Delta_{j}^{2},\ldots,\Delta_{j}^{n})$ is the population median of the paired differences for feature $j$. The test operates by ranking the absolute values of the non-zero paired differences, then comparing the sum of ranks associated with positive differences against those associated with negative differences. This non-parametric approach is robust to outliers and skewed distributions, making it well-suited for stylometric features that may not follow a Gaussian distribution. As we perform one test per feature, we control for multiple comparisons using the Benjamini-Hochberg false discovery rate (FDR) correction Benjamini and Hochberg (1995) across all features. We report a significance indicator in our tables.
Finally, features are ranked by the absolute value of their paired Cohen’s $d$, and the top-ranked features are reported along with their FDR-corrected significance status. This ranking highlights the stylometric dimensions along which AI-generated persuasive text diverges most strongly from its human-written counterpart.